Need help with this question. Thank you in advance!

You are a farmer. You have two plots of land. One is on a hill and one is in a valley. Each is ten acres. On each plot, you can grow either corn or wheat. Each acre in the valley will produce either 10 bushels of wheat or 5 bushels of corn. Each acre of the hill-top plot will produce either 6 bushels of wheat or 4 bushels of corn.

If the price of corn is $10 and the price of wheat is $8, what would you plant and where in order to maximize revenue? How much would your crops be worth?
Answer #1

Revenue through each acre of valley if wheat is produced = 10*8=80.

Revenue through each acre of valley if corn is produced=5*10=50

Therefore wheat would be produced in the valley as revenue from wheat is higher in the valley.

On hill top, revenue from one acre if wheat is produced =6*8=48

On.hill top, revenue from one acre if corn is produced = 4*10= 40.

Therefore producing wheat on hill top will generate more revenue than corn, so wheat would be produced on hill top too.

Total worth of crops from valley = 80*10 dollars = 800 dollars

Total worth of crops from.hill.top = 48*10= 480 dollars

Therefore total worth = 800+480 = 1280 dollars